When 1956 J of heat energy is added to 44.1 g of hexane, C6H14, the temperature increases by 19.6*C.

calculate the molar heat capacity of C6H14


1
Expert's answer
2021-04-09T01:46:01-0400

Molar mass of "C_6H_{14}" is "86.176gmol^{-1}"

number of moles, "n=\frac{mass}{molar \space mass}=\frac{44.1g}{86.176gmol^{-1}}=0.512mol"

molar heat capacity of "C_6H_{14}, C=\frac{q}{n\Delta T}=\frac{1956}{(0.512)(19.6)}=194.91Jmol^{-1o}C^{-1}"